# Confronting recent light compact star observations with color-flavor locked quark matter

### v1

@@ 4.2 Hybrid stars
Notably, the simultaneous reconciliation of both HESS J1731-347 and PSR J1231-1431 (in Case I) measurements would be rather difficult assuming a purely hadronic EOS. In particular, the hadronic model should be rather soft, at low densities, to support the low radius associated with HESS J1731-347 in the sub-solar mass region, and then it should rapidly stiffen to achieve the higher radius of PSR J1231-1431 at slightly larger masses. Interestingly, the low-density domain of the nuclear EOS can be effectively constrained through parity-violating electron scattering experiments, which aim to measure the neutron skin thickness of different nuclei. In particular, the PREX-II collaboration provided a measurement for the neutron skin thickness of lead, which pointed to stiff EOS behavior at low densities (74, 75). Then, the subsequent CREX experiment extracted a puzzling value for the neutron skin thickness of calcium (76), supporting softer models. More precisely, the simultaneous explanation of both CREX and PREX-II was not possible with the use of traditional energy density functionals (77, 78, 79, 80, 81). However, recent attempts have proposed sophisticated modifications to the Lagrangians which are used to describe nuclear matter, and they achieved an explanation of both experimental values (82, 83). However, the predicted EOSs are characterized as extremely stiff. Notably, the latter issue was at some extent resolved in Ref. (84).
@@
Figure 3: Mass-radius diagrams for hybrid EOSs combining the Ska hadronic EOS and the CFL MIT bag model. The gray contour regions denote mass and radius measurements related to PSR J0740+6620 (37), PSR J0030+0451 (40), HESS J1731-347 (7), PSR J0437-4715 (36) , PSR J1231-1411 (31), XTE J1814-338 (63) and GW170817 (72). The solid (dashed) contours correspond to $2\sigma\ (1\sigma)$ confidence. Panel (a) shows the results for $a_{4}=1$. Panel (b) shows the results for $a_{4}=0.7$. Panel (c) includes the results for $a_{4}=1$ including the contribution of a vector interaction among quarks with coupling $G_{v}=0.15\ \mathrm{fm^{2}}$.
@@
Considering that current experimental constraints might point to stiff behavior for the hadronic EOS at low density, in the present study, we are going to employ a nuclear model which is sufficiently stiff, so that it crosses the PSR J1231-1411 contour (Case I), and, therefore, potentially incompatible to the HESS J1731-347 constraints. In particular, we are going to use a widely employed Skyrme model, namely Ska (85, 86, 87, 88, 89). Thus, in our attempt for explaining all current astronomical measurements, by considering a first-order phase transition to CFL quark matter, we will rely on the hybrid branch to account for the existence of the central compact object in the HESS J1731-347 remnant. It is important to comment that, the employed nuclear model is of nucleonic composition. In principle, as density increases, hyperonic degrees of freedom may appear, altering the properties of the EOS. Given the uncertainties related to hyperon-hyperon and hyperon-nucleon interactions, in the present study we work under a simplified framework that neglects their existence.
@@
In Fig. 3(a), we depicted the mass-radius dependence for hybrid EOSs constructed by varying the values of $\Delta$ and $B$ (fixing $a_{4}=1$). Notably, the pairs of $\Delta$ and $B$ are selected so that the phase transition density remains fixed, to allow for the hadronic branch to account for PSR J1231-1411. As one can observe, the hybrid branch crosses the HESS J1731-347 inferred region. However, the EOS softening that is induced due to the phase transition plays a critical role on the determination of the resulting maximum mass. As a consequence, the hybrid branch never reaches the $2M_{\odot}$ and, therefore, it fails to reproduce one of the most robust constraints derived by compact object observations. In Fig. 3(b) we attempted to resolve this issue by lowering the value of the parameter $a_{4}$. The motivation to do that relies on the knowledge that reducing $a_{4}$ reduces the value of pressure for a given baryon chemical potential and, as a consequence, for the same parameter values of $B$ and $\Delta$ shifts the onset of the phase transition to higher densities. Therefore, to recover the desired phase transition onset (i.e., when the hadronic branch crosses the PSR J1231-1411 contour) we need to reduce the value of the bag constant (90), leading to EOS stiffening. While the constructed EOSs achieve higher masses (still not the conservative $2M_{\odot}$ limit), the induced stiffening does not allow for an explanation of the HESS J1731-347 measurement. Finally, we tried to stiffen the quark EOS by adding an ad hoc vector interaction between quarks. We refer to this attempt as ad hoc since the resulting models do not follow from an initial unified description, but are produced by simply adding a vector mean-field term in the CFL MIT bag model expressed in the canonical ensemble (following Refs. (91, 92, 93), the strength of the vector repulsion is quantified via the parameter $G_{v}=(g/m_{v})^{2}$, where $g$ is the coupling constant and $m_{v}$ the mass of the mediating boson). For more details on the inclusion of this vector term the reader is referred to Ref. (91). In any case, the results depicted in Fig. 3(c) indicate that such an approach also does not suffice to reconcile all of the considered astronomical constraints. However, while the CFL MIT bag model might fail in reproducing all measurements (when assuming a phase transition), we can not exclude that a different, comprehensive analysis (e.g., via a relativistic mean-field type model), that includes such a stiffening term, may enable the simultaneous explanation of all observations.
@@
Figure 4: (a) Mass-radius diagrams for hybrid EOSs combining the Ska and the CFL bag model with $a_{4}=1$. Parameters $B$ and $\Delta$ are chosen so that the phase transition occurs when the hadronic branch reaches $2M_{\odot}$. Each curve is drawn until the last slow stable configuration. The observational data are the same as in Fig. 3. (b) Comparing the non CFL contribution to the quark EOSs to pQCD results (97, 98). The ‘x’ marks denote the last slow stable configuration.
@@
Interestingly, we can attempt to reconcile all of the considered astronomical constraints by allowing the hadronic branch to reach the two solar masses and then induce an extremely strong phase transition to make the mass-radius diagram drop to cross the HESS J1731-347 contour. Notably, hybrid stars may be stable even at a descending branch of a mass-radius diagram, assuming that the phase conversion is sufficiently slow (94). The characterization slow refers to the magnitude of the phase conversion timescale $\tau_{\mathrm{conv}}$, compared to the period $\tau_{\mathrm{osc}}$ of radial oscillations that the stellar configurations may undergo. In particular, if $\tau_{\mathrm{conv}}>>\tau_{\mathrm{osc}}$, then hybrid stars could be stable even at regions where the well-known turning point criterion would rule them out (94, 95). The idea of explaining the HESS J1731-347 constraints as slow stable hybrid star is not new. More precisely, Mariani et al. (18) established the plausibility of this scenario in their recent work. In the present study, we aim to examine if the CFL MIT bag model can lead to such an explanation. Specifically, the authors of Ref. (18) considered a constant speed of sound parametrization where there are three free parameters, namely the transition pressure, the speed of sound and the energy density jump. As a consequence, one can construct a sufficiently stiff EOS, by controlling the speed of sound, which would be characterized by a large energy density jump at any chosen transition density. In contrast, in the current framework we have two free parameters (we are keeping the value of $a_{4}=0.8$ fixed to account for pQCD constraints; see below) and their interplay has to self-consistently provide an appropriate transition density (so that the hadronic branch reaches the $2M_{\odot}$), and a large enough energy density discontinuity to trigger an almost vertical drop of the mass-radius curve. Therefore, it is particularly interesting to examine which pairs of the phenomenological parameters $B$ and $\Delta$ could lead to a simultaneous explanation of all current astronomical constraints.
@@
In Fig. 4(a), we depict the mass-radius dependence for hybrid EOSs constructed by varying the color superconducting gap in the range [0,600] MeV. The bag constant was derived by considering that the phase transition occurs when the hadronic mass reaches the $2M_{\odot}$. Notably, all of the hybrid branches, in Fig. 4(a), are plotted until the last slow stable configuration or terminal configuration as named in Ref. (18). As one can observe, the reconciliation of the HESS J1731-347 constraints is possible, but it would require extremely large values for both $B$ and $\Delta$ (see the legend of Fig. 4), at least compared to those found usually in the literature. This is, however, somewhat expected as the majority of studies focuses on phase transition of moderate energy density jumps and low to moderate transition densities. Hence, the extreme conditions required for the explanation of HESS J1731-347 require extreme values for the phenomenological parameters of CFL quark matter. In any case, it is worth noting that recent and robust constraints have been imposed on the color superconduncting gap in Ref. (96). While a value of around $\sim$ 400 MeV is potentially compatible to their results, larger values may be disfavored (see Fig. 2 in Ref. (96)).
@@
Notably, the central baryon chemical potential range that is being considered, in order to achieve the reconciliation of the HESS J1731-347 constraints, reaches such high values that it may cross the regime where the results of pQCD, for the EOS of strongly interacting quark matter, are potentially reliable. Typically, pQCD is considered to be credible at densities around $\sim 40n_{0}$ (where $n_{0}$ is the nuclear saturation density), which correspond to values of baryon chemical potential that most likely exceed 2 GeV (96, 97, 98, 99, 100, 2, 101). For that matter, we wanted to test whether the constructed EOSs are compatible to pQCD results (for values beyond 2 GeV), at least for values of baryon chemical potential equal to those found at the center of the last slow stable configurations. In Fig. 4(b), we plotted the quark EOSs, compared to $\mathcal{O}(a_{s}^{2})$ results of pQCD (97) (employing the fit to the numerical data provided by Fraga et al. (98)). Notably, each curve depicts the pressure (divided by the Stefan-Boltzmann pressure $P_{SB}=3(\mu_{B}/3)^{4}/(4\pi^{2}\hbar^{3})$) as a function of the chemical potential, neglecting the contribution of color superconductivity (as such a contribution is not encapsulated in the considered pQCD calculations (98)). Finally, the ’x’ appearing on each curve denotes the terminal configuration for each EOS. As is evident, for values of $\Delta$ up to $400$ MeV the derived EOSs are compatible to the results of pQCD. However, this is not the case for larger $\Delta$ values. The latter is of utmost importance as it places strong constraints on the possible range of parameters that the enable explanation of the HESS J1731-347 constraints in a slow phase conversion scenario.
@@
A final remark is appropriate regarding the fact that while the existence of slow stable hybrid stars in theoretically intriguing it is not clear how such objects are born. In that sense, future work on possible formation scenarios of such objects would be of utmost importance. Some recent progress on the astrophysical paths that may lead to the existence of twin stars (but for a hybrid branch respecting the turning point criterion) has been made in the recent work of Ref. (56). In that direction, we expect that future research will hopefully shed light on the possible existence of slow stable hybrid star branches.
@@ 5 Conclusions
When considering the framework of hybrid stars, we found that although a hybrid branch originating in the PSR J1231-1411 mass-radius region (as provided by Ref. (31)) can also accommodate the HESS J1731-347 measurement, the resulting maximum masses remain well below the well-established threshold of $2M_{\odot}$. While altering the pQCD related parameter $a_{4}$ or incorporating a repulsive vector interaction increases the predicted maximum mass, the induced stiffening does not allow for the explanation of the HESS J1731-347 constraints (assuming a stiff hadronic EOS). However, the situation changes if the CCO in the HESS J1731-347 remnant is interpreted as a slow stable hybrid star. In this scenario, we choose a sufficiently strong phase transition at high central pressures, requiring high values for both $\Delta$ and $B$, to produce a hybrid branch that initiates at $2M_{\odot}$ and drops thereafter, while maintaining stability. We concluded that such a hybrid EOS can be compatible with all of the aforementioned measurements, while also satisfying the demands of pQCD at extremely high density.

### v2

@@ 4.2 Hybrid stars
Notably, the simultaneous reconciliation of both HESS J1731-347 and PSR J1231-1431 (in Case I) measurements would be rather difficult assuming a purely hadronic EOS. In particular, the hadronic model should be rather soft, at low densities, to support the low radius associated with HESS J1731-347 in the sub-solar mass region, and then it should rapidly stiffen to achieve the higher radius of PSR J1231-1431 at slightly larger masses. Interestingly, the low-density domain of the nuclear EOS can be effectively constrained through parity-violating electron scattering experiments, which aim to measure the neutron skin thickness of different nuclei. In particular, the PREX-II collaboration provided a measurement for the neutron skin thickness of lead, which pointed to stiff EOS behavior at low densities (86, 87). Then, the subsequent CREX experiment extracted a puzzling value for the neutron skin thickness of calcium (88), supporting softer models. More precisely, the simultaneous explanation of both CREX and PREX-II was not possible with the use of traditional energy density functionals (89, 90, 91, 92, 93). However, recent attempts have proposed sophisticated modifications to the Lagrangians which are used to describe nuclear matter, and they achieved an explanation of both experimental values (94, 95). Nonetheless, the predicted EOSs are characterized as extremely stiff. Notably, the latter issue was at some extent resolved in Ref. (96).
@@
Figure 3: Mass-radius diagrams for hybrid EOSs combining the Ska hadronic EOS and the CFL MIT bag model. The gray contour regions denote mass and radius measurements related to PSR J0740+6620 (37), PSR J0030+0451 (40), HESS J1731-347 (7), PSR J0437-4715 (36) , PSR J1231-1411 (31), XTE J1814-338 (70) and GW170817 (84). The solid (dashed) contours correspond to $2\sigma\ (1\sigma)$ confidence. (a) shows the results for $a_{4}=1$. (b) shows the results for $a_{4}=0.7$. (c) includes the results for $a_{4}=1$ including the contribution of a vector interaction among quarks with coupling $G_{v}=0.15\ \mathrm{fm^{2}}$.
@@
Considering that current experimental constraints might point to stiff behavior for the hadronic EOS at low density, in the present study, we are going to employ a nuclear model which is sufficiently stiff, so that it crosses the PSR J1231-1411 contour (Case I), and, therefore, potentially incompatible to the HESS J1731-347 constraints. In particular, we are going to use a widely employed Skyrme model, namely Ska (97, 98, 99, 100, 101). Thus, in our attempt for explaining all current astronomical measurements, by considering a first-order phase transition to CFL quark matter, we will rely on the hybrid branch to account for the existence of the central compact object in the HESS J1731-347 remnant. It is important to comment that, the employed nuclear model is of nucleonic composition. In principle, as density increases, hyperonic degrees of freedom may appear, altering the properties of the EOS. Given the uncertainties related to hyperon-hyperon and hyperon-nucleon interactions, in the present study we work under a simplified framework that neglects their existence.
@@
In Fig. 3(a), we depicted the mass-radius dependence for hybrid EOSs constructed by varying the values of $\Delta$ and $B$ (fixing $a_{4}=1$). Notably, the pairs of $\Delta$ and $B$ are selected so that the phase transition density remains fixed, to allow for the hadronic branch to account for PSR J1231-1411. As one can observe, the hybrid branch crosses the HESS J1731-347 inferred region. However, the EOS softening that is induced due to the phase transition plays a critical role on the determination of the resulting maximum mass. As a consequence, the hybrid branch never reaches the $2M_{\odot}$ and, therefore, it fails to reproduce one of the most robust constraints derived by compact object observations. In Fig. 3(b) we attempted to resolve this issue by lowering the value of the parameter $a_{4}$. The motivation to do that relies on the knowledge that reducing $a_{4}$ reduces the value of pressure for a given baryon chemical potential and, as a consequence, for the same parameter values of $B$ and $\Delta$, shifts the onset of the phase transition to higher densities. Therefore, to recover the desired phase transition onset (i.e., when the hadronic branch crosses the PSR J1231-1411 contour) we need to reduce the value of the bag constant (102), leading to EOS stiffening. While the constructed EOSs achieve higher masses (still not the conservative $2M_{\odot}$ limit), the induced stiffening does not allow for an explanation of the HESS J1731-347 measurement. Finally, we tried to stiffen the quark EOS by adding an ad hoc vector interaction between quarks. We refer to this attempt as ad hoc since the resulting models do not follow from an initial unified description, but are produced by simply adding a vector mean-field term in the CFL MIT bag model expressed in the canonical ensemble (following Refs. (103, 104, 105), the strength of the vector repulsion is quantified via the parameter $G_{v}=(g/m_{v})^{2}$, where $g$ is the coupling constant and $m_{v}$ the mass of the mediating boson). For more details on the inclusion of this vector term the reader is referred to Ref. (103). In any case, the results depicted in Fig. 3(c) indicate that such an approach also does not suffice to reconcile all of the considered astronomical constraints. However, while the CFL MIT bag model might fail in reproducing all measurements (when assuming a phase transition), we can not exclude that a different, comprehensive analysis (e.g., via a relativistic mean-field type model), that includes such a stiffening term, may enable the simultaneous explanation of all observations.
@@
Figure 4: (a) Mass-radius diagrams for hybrid EOSs combining the Ska and the CFL bag model with $a_{4}=1$. Parameters $B$ and $\Delta$ are chosen so that the phase transition occurs when the hadronic branch reaches $2M_{\odot}$. Each curve is drawn until the last slow stable configuration. The observational data are the same as in Fig. 3. (b) Comparing the non CFL contribution to the quark EOSs to pQCD results (109, 110). The ‘x’ marks denote the last slow stable configuration.
@@
Interestingly, we can attempt to reconcile all of the considered astronomical constraints by allowing the hadronic branch to reach the two solar masses and then induce an extremely strong phase transition to make the mass-radius diagram drop to cross the HESS J1731-347 contour. Notably, hybrid stars may be stable even at a descending branch of a mass-radius diagram, assuming that the phase conversion is sufficiently slow (106). The characterization slow refers to the magnitude of the phase conversion timescale $\tau_{\mathrm{conv}}$, compared to the period $\tau_{\mathrm{osc}}$ of radial oscillations that the stellar configurations may undergo. In particular, if $\tau_{\mathrm{conv}}>>\tau_{\mathrm{osc}}$, then hybrid stars could be stable even at regions where the well-known turning point criterion would rule them out (106, 107). The idea of explaining the HESS J1731-347 CCO as a slow stable hybrid star is not new. More precisely, Mariani et al. (18) established the plausibility of this scenario in their recent work. In the present study, we aim to examine if the CFL MIT bag model can lead to such an explanation. Specifically, the authors of Ref. (18) considered a constant speed of sound parametrization where there are three free parameters, namely the transition pressure, the speed of sound and the energy density jump. As a consequence, one can construct a sufficiently stiff EOS, by controlling the speed of sound, which would be characterized by a large energy density jump at any chosen transition density. In contrast, in the current framework we have two free parameters (we are keeping the value of $a_{4}=0.8$ fixed to account for pQCD constraints; see below) and their interplay has to self-consistently provide an appropriate transition density (so that the hadronic branch reaches the $2M_{\odot}$), and a large enough energy density discontinuity to trigger an almost vertical drop of the mass-radius curve. Therefore, it is particularly interesting to examine which pairs of the phenomenological parameters $B$ and $\Delta$ could lead to a simultaneous explanation of all current astronomical constraints.
@@
In Fig. 4(a), we depict the mass-radius dependence for hybrid EOSs constructed by varying the color superconducting gap in the range [0,600] MeV. The bag constant was derived by considering that the phase transition occurs when the hadronic mass reaches the $2M_{\odot}$. Notably, all of the hybrid branches, in Fig. 4(a), are plotted until the last slow stable configuration or terminal configuration as named in Ref. (18). As one can observe, the reconciliation of the HESS J1731-347 constraints is possible, but it would require extremely large values for both $B$ and $\Delta$ (see the legend of Fig. 4), at least compared to those found usually in the literature. This is, however, somewhat expected as the majority of studies focuses on phase transitions of moderate energy density jumps and low to moderate transition densities. Hence, the extreme conditions required for the explanation of HESS J1731-347 require extreme values for the phenomenological parameters of CFL quark matter. In any case, it is worth noting that recent and robust constraints have been imposed on the color superconduncting gap in Ref. (108). While a value of around $\sim$ 400 MeV is potentially compatible to their results, larger values may be disfavored (see Fig. 2 in Ref. (108)).
@@
Notably, the central baryon chemical potential range that is being considered, in order to achieve the reconciliation of the HESS J1731-347 constraints, reaches such high values that it may cross the regime where the results of pQCD, for the EOS of strongly interacting quark matter, are potentially reliable. Typically, pQCD is considered to be credible at densities around $\sim 40n_{0}$ (where $n_{0}$ is the nuclear saturation density), which correspond to values of baryon chemical potential that most likely exceed 2 GeV (108, 109, 110, 111, 112, 2, 113). For that matter, we wanted to test whether the constructed EOSs are compatible to pQCD results (for values beyond 2 GeV), at least for values of baryon chemical potential equal to those found at the center of the last slow stable configurations. In Fig. 4(b), we plotted the quark EOSs, compared to $\mathcal{O}(a_{s}^{2})$ results of pQCD (109) (employing the fit to the numerical data provided by Fraga et al. (110)). Notably, each curve depicts the pressure (divided by the Stefan-Boltzmann pressure $P_{SB}=3(\mu_{B}/3)^{4}/(4\pi^{2}\hbar^{3})$) as a function of the chemical potential, neglecting the contribution of color superconductivity (as such a contribution is not encapsulated in the considered pQCD calculations (110)). Finally, the ’x’ appearing on each curve denotes the terminal configuration for each EOS. As is evident, for values of $\Delta$ up to $400$ MeV the derived EOSs are compatible to the results of pQCD. However, this is not the case for larger $\Delta$ values. The latter is of utmost importance as it places strong constraints on the possible range of parameters that may enable the explanation of the HESS J1731-347 constraints in a slow phase conversion scenario.
@@
A final remark is appropriate regarding the fact that while the existence of slow stable hybrid stars is theoretically intriguing it is not clear how such objects are born. In that sense, future work on possible formation scenarios of such objects would be of utmost importance. Some recent progress on the astrophysical paths that may lead to the existence of twin stars (but for a hybrid branch respecting the turning point criterion) has been made in the work of Ref. (56). In that direction, we expect that future research will hopefully shed light on the possible existence of slow stable hybrid star branches.
@@ 5 Conclusions
Notably, in the case of absolutely stable SQM, we worked under the assumption that only one state of matter may appear in the stellar interior. However, it is worth mentioning that a new exotic scenario has been reported in the literature suggesting the possible reappearance of hadrons, at large densities, even if SQM represents the true ground state of matter (114, 115, 116). Thus, an interesting direction for future work would be to investigate how such a consideration might alter the results reported in the present study.
When considering the framework of hybrid stars, we found that although a hybrid branch originating in the PSR J1231-1411 mass-radius region (as provided by Ref. (31)) can also accommodate the HESS J1731-347 measurement, the resulting maximum masses remain well below the well-established threshold of $2M_{\odot}$. While altering the pQCD related parameter $a_{4}$ or incorporating a repulsive vector interaction increases the predicted maximum mass, the induced stiffening does not allow for the explanation of the HESS J1731-347 constraints (assuming a stiff hadronic EOS). However, the situation changes if the CCO in the HESS J1731-347 remnant is interpreted as a slow stable hybrid star. In this scenario, we adopted a sufficiently strong phase transition at high central pressures, requiring high values for both $\Delta$ and $B$, to produce a hybrid branch that initiates at $2M_{\odot}$ and drops thereafter, while maintaining stability. We concluded that such a hybrid EOS can be compatible with all of the aforementioned measurements, while also satisfying the demands of pQCD at extremely high density.